How much charge passes through a wire in 2.5 minutes if the current in the wire is 1.6 A? 


1
Expert's answer
2021-12-20T10:27:28-0500

"Q = I \cdot t \\
I = 1.6 \: A \: \\
t = 2.5 \: min = 2.5 \cdot 60 = 150 \: s \\
Q = I \cdot t = 1.6 \cdot 150 = 240 \: Coulomb \\
Answer: Q = 240 \: C"